If x 75 = y 45 = z 15 = 0 , then which of the statement is true
(a) x + y = 2z (b) x + y = 3z (c) x – y = 2z (d) x – y = 3z

Grade:9

1 Answers

Ravi
9 years ago
If you mean x/75=y/45=z/15

Then letx/75=y/45=z/15 =k

So.x=75k y=45k z=15k

Pu the value in the options to get the correct option satisfying the equality
.....................................................................................................................................